OpenStudy (anonymous):

The volume of a cylinder is 81 ft3. The volume of a cube is 42 ft3. What is the difference between the volume of a cone that fits exactly inside the cylinder and the volume of a pyramid that fits exactly inside the cube? Provide an explanation for your answer.

OpenStudy (anonymous):

do you know what the volume formula is

OpenStudy (anonymous):

Hi sorry

OpenStudy (anonymous):

It's v=pi*r^2*h

OpenStudy (anonymous):

\[V=l*w*h\]

OpenStudy (anonymous):

Volume of a cylinder

OpenStudy (anonymous):

hey srry ur right

OpenStudy (anonymous):

Okay so what do I do from here

OpenStudy (anonymous):

you check your facebook for new replies :DD

OpenStudy (anonymous):

I don't have facebook sooo

OpenStudy (anonymous):

@Abmon98 dude

OpenStudy (anonymous):

so plug in the numbers for both of the equations and then compare them... and tell me if i did something wrong

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

oh you don't? It's hard to believe a teenage who doesn't have FB

OpenStudy (anonymous):

i dont have one

OpenStudy (anonymous):

Nah I have a twitter, follow me on there

OpenStudy (anonymous):

i have a twitter

OpenStudy (anonymous):

Mine is swatttts

OpenStudy (anonymous):

i love wrestling can someone give me a medal please

OpenStudy (anonymous):

Now tell me the answer ._.

OpenStudy (anonymous):

SOMEONE

OpenStudy (anonymous):

just copy and pace the formulas and stuff i said and do you take flvs

OpenStudy (anonymous):

Omg so confusing no

OpenStudy (anonymous):

its really easy actually

OpenStudy (anonymous):

@whpalmer4 start tagging people with a high smartscore so you can get the answer @nurali will give answers and will get them correct for you

OpenStudy (anonymous):

so fan her

OpenStudy (anonymous):

@Abmon98 Is awesome

OpenStudy (abmon98):

V=pir^2*h cylinder v=1/3pir^2h cone v=l*w*h V=1/3b^2h 1/3 of the volume can fit a whole shape of a cone

OpenStudy (anonymous):

im not:(

OpenStudy (abmon98):

and its the same thing to the pyramid and cube he whole shape(pyriamid can fit in only 1/3 of the cube

OpenStudy (abmon98):

Thank you @savvvvannahh :D